# A short proof of generalized Conway–Gordon–Sachs theorem

### v1

@@ 1. Introduction
An embedding $f$ of a finite graph $G$ into the $3$-sphere is called a spatial embedding of $G$, and $f(G)$ is called a spatial graph of $G$. We call a subgraph of $G$ homeomorphic to the circle a cycle of $G$, and a cycle containing exactly $p$ vertices a $p$-cycle. We denote the set of all $p$-cycles of $G$ by $\Gamma_{p}(G)$. Additionally, we also denote the set of all disjoint pairs of cycles of $G$ consisting of a $p$-cycle and a $q$-cycle by $\Gamma_{p,q}(G)$. In general, for a closed $1$-manifold $\lambda$ in $G$, the image $f(\lambda)$ is no other than a knot or a link contained in $f(G)$. In particular, if $\lambda$ belongs to $\Gamma_{p,q}(G)$, then we call $f(\lambda)$ a $(p,q)$-link of $f(G)$. Moreover, if $\lambda$ contains all vertices of $G$, then we also call $f(\lambda)$ a Hamiltonian link of $f(G)$.
@@
Kazakov–Korablev’s result above can be obtained immediately by taking modulo $2$ reduction on both sides of (1.2). Furthermore, since the right side of (1.2) is not zero, it is also clear that for any $p$ and $q$ satisfying $p+q=n$, there always exists a nonsplitable Hamiltonian $(p,q)$-link of $f(K_{n})$. This is also a generalization of Vesnin–Litvintseva’s result [8] that $f(K_{n})$ always contain a nonsplittable Hamiltonian link. For additional formulas on evaluating the sum of linking numbers over Hamiltonian links derived from (1.1), see [4]. The proof of Theorem 1-1 given in [4] was done using tricky induction procedures and was lengthy and tedious. Our purpose in this paper is to simplify and shorten the proof of Theorem 1-1, thereby making it more accessible to a wider reader. The key to this simplification is Lemma 2-2, which reveals a particular homological property concerning the linking of a knot with a spatial complete graph. The proof of Theorem 1-1 presented here is a reorganized version of the one originally introduced by the author in the Japanese textbook [6] several years ago.
@@ 2. New proof of Theorem 1-1
For a $k$-cycle $\gamma$ of $K_{n}$, we denote the subgraph of $K_{n}$ obtained by removing all $k$ vertices of $\gamma$ and all edges incident to them from $K_{n}$ by $G_{\gamma}$. Note that this subgraph is isomorphic to $K_{n-k}$. Then we have
@@
Let $H_{\gamma^{\prime}}^{i}\ (i=1,2,\ldots,l,\ l=\binom{n-3}{p})$ be the collection of all subgraphs of $G_{\gamma^{\prime}}$ that are isomorphic to $K_{p}$. Then the sets of $p$-cycles $\Gamma_{p}(H_{\gamma^{\prime}}^{i})$ are mutually disjoint, and their direct sum equals $\Gamma_{p}(G_{\gamma^{\prime}})$. Then by Corollary 2-3, we have
@@
Note that each $3$-cycle $\gamma$ of $G_{\gamma^{\prime}}$ is shared by exactly $\binom{n-6}{p-3}$ subgraphs $H_{\gamma^{\prime}}^{i}$’s. This implies that

### v3

# A short proof of the generalized Conway–Gordon–Sachs theorem

@@ 1. Introduction
An embedding $f$ of a finite graph $G$ into the $3$-sphere is called a spatial embedding of $G$, and $f(G)$ is called a spatial graph of $G$. We call a subgraph of $G$ homeomorphic to the circle a cycle of $G$, and a cycle containing exactly $p$ vertices a $p$-cycle. We denote the set of all $p$-cycles of $G$ by $\Gamma_{p}(G)$. Additionally, we also denote the set of all disjoint pairs of cycles of $G$ consisting of a $p$-cycle and a $q$-cycle by $\Gamma_{p,q}(G)$. In general, for a closed $1$-manifold $\lambda$ in $G$, the image $f(\lambda)$ is a knot or a link contained in $f(G)$. In particular, if $\lambda$ belongs to $\Gamma_{p,q}(G)$, then we call $f(\lambda)$ a $(p,q)$-link of $f(G)$. Moreover, if $\lambda$ contains all vertices of $G$, then we also call $f(\lambda)$ a Hamiltonian link of $f(G)$.
@@
Kazakov–Korablev’s result above can be obtained immediately by taking the modulo $2$ reduction on both sides of (1.2). Since the right side of (1.2) is not zero, it is also clear that for any $p$ and $q$ satisfying $p+q=n$, there always exists a nonsplitable Hamiltonian $(p,q)$-link of $f(K_{n})$. This is also a generalization of Vesnin–Litvintseva’s result [8] that $f(K_{n})$ always contain a nonsplittable Hamiltonian link. Furthermore, it has also been shown in [4] that for $n\geq 6$, the following congruence modulo $2\cdot(n-5)!$ can be obtained from (1.2):
[…]
For additional formulas on evaluating the sum of linking numbers over Hamiltonian links derived from (1.1), see [4]. The proof of Theorem 1-1 given in [4] was done using tricky induction procedures and was lengthy and tedious. Our purpose in this paper is to simplify and shorten the proof of Theorem 1-1, thereby making it more accessible to a wider audience. The key to this simplification is Lemma 2-2, which reveals a particular homological property concerning the linking of a knot with a spatial complete graph. The proof of Theorem 1-1 presented here is a reorganized version of the one originally introduced by the author in the Japanese textbook [6] several years ago.
@@ 2. New proof of Theorem 1-1
For a $k$-cycle $\gamma$ of $K_{n}$, let $G_{\gamma}$ denote the subgraph of $K_{n}$ obtained by removing the $k$ vertices of $\gamma$ and their edges. Note that this subgraph is isomorphic to $K_{n-k}$. Then we have
@@
Let $H_{\gamma^{\prime}}^{i}\ (i=1,2,\ldots,l,\ l=\binom{n-3}{p})$ be the collection of all subgraphs of $G_{\gamma^{\prime}}$ that are isomorphic to $K_{p}$. Then the sets of $p$-cycles $\Gamma_{p}(H_{\gamma^{\prime}}^{i})$ are mutually disjoint, and their union is $\Gamma_{p}(G_{\gamma^{\prime}})$. Then by Corollary 2-3, we have
@@
Note that each $3$-cycle $\gamma$ of $G_{\gamma^{\prime}}$ is shared by exactly $\binom{n-6}{p-3}$ of the $H_{\gamma^{\prime}}^{i}$ subgraphs. This implies that